Please show explanation. I am really strugglPlease show explanation. I am really struggling with this part of circuits. I will give great review and rating.
Answer #1

Applying Kirchoffs Law

Here, As I1 = I2 + I3

=>   18 + 3 = (I2 + I3) * 20.5 + 10.5 * I3

Also, 18 + 12 - 24 = (I2 + I3) * 20.5 + 7 * I2

=> Solving for I1 , I2 , I3

=>   I2 = - 0.5656 A

=> I3 = 1.0514 A

=> I1 =   1.0514 - 0.5656

          =   0.4858 A